A particle falling under gravity describes 80ft in certain second. How long does it take to describe the next 112ft?

Grade:7

1 Answers

Nirmal Singh.
10 years ago
let us consider the velocity of particle when it completes 80 ft is v and this velocity will be initial velocity for the 112 ft height. let us consider it takes t time.
g = 32 ft/s^2
v^2 = u^2 + 2gh
v^2 = 0^2+ 2* 32* 80
v = sqrt (64*80)
v = 8*4*sqrt(5)
v = 71.55 m/s
now h = vt+0.5*g*t^2
112 = 71.55t+0.5*32*t^2
On solving
t = -b+(-)sqrt(b^2-4ac) / 2a
t = 1.2 sec
Thanks & Regards,
Nirmal Singh
Askiitians faculty